OpenStudy (anonymous):

Find the component form and magnitude of the indicated vector: Given that P=(-10,15) and Q=(-8,8), find the component form and magnitude of the vector sqrt5 QP

OpenStudy (amistre64):

i assume QP means from Q to P therefore the vector is end - start; P-Q

OpenStudy (amistre64):

sqrt(5) means we scale that result by a factor of sqrt(5) ; then take the length of it

OpenStudy (anonymous):

What?

OpenStudy (amistre64):

given that these are all terms of vector mathing, I have to wonder just what you means by "what?"

OpenStudy (anonymous):



OpenStudy (amistre64):

yep, thats the condition that I have presented

OpenStudy (anonymous):

so what exactly do i do with it?

OpenStudy (amistre64):

... subtract Q from P. what do we get? P-Q = ?

OpenStudy (anonymous):

<2,-7>

OpenStudy (amistre64):

good, this is our vector from Q to P scale that by sqrt(5) by multiplying the components by sqrt(5)

OpenStudy (amistre64):

15-8 = 7 tho; so drop the negative there :)

OpenStudy (anonymous):

isnt it q -p though?

OpenStudy (amistre64):

-10--8 = -2 15-8 = 7 ; so the direction is imperitive. <-2,7>

OpenStudy (amistre64):

end - start = direction

OpenStudy (amistre64):

we end in P and start at Q

OpenStudy (anonymous):

\[2\sqrt{5} , 7\sqrt{5}\] ?

OpenStudy (amistre64):

getting better, lets clean it up and make sure we know how we got there :) <-2sqrt(5) , 7sqrt(5)>

OpenStudy (amistre64):

P(-10,15) -Q( -8, 8) ----------- < -2, 7> *sqrt(5)

OpenStudy (amistre64):

then the magnitude is to square the parts, add them together, and sqrt it

OpenStudy (amistre64):

sqrt(4*5 + 49*5)

OpenStudy (anonymous):

That's what I had gotten the first time and my teacher marked it off on my test...guess I'll go ask her if she got something wrong. Thanks so much!

OpenStudy (amistre64):

your welcome; make sure youve simplified the sqrt(..) as far as possible

OpenStudy (amistre64):

not that 5 and 53 have any perfect squares to pull out

OpenStudy (anonymous):

Thank you so much!

OpenStudy (amistre64):

youre welcome.